if a car takes 20 minutes to travel 15km how fast is the car going ?
Answer by mananth(16946)

if a car takes 20 minutes to travel 15km how fast is the car going ?
..
since distance is in km
time has to be changed to hours
20 minutes = 20/60 = 1/3 hours
..
speed = distance / time
speed = 15 / (1/3)
speed = 15*3
= 45 km / hour
..
Your feedback is important to improve the solution